# A Million Three-body Binaries Caught by Gaia

### v2

@@ 3 Mock Three-body Binary Population versus Gaia Wide Binaries
Figure 1 displays the total number of solar neighborhood 3BBs we predict across all our models, alongside the wide binary ($\langle s\rangle>10^{3}\rm AU$) tally. The “tidal isolation” parameter, $\eta_{H}$ (§ A.1), is a nuisance parameter quantifying how isolated a three-body encounter is from tidal perturbations applied by its host cluster. Smaller $\eta_{H}$ demands increasingly isolated three-body encounters, with $\eta_{H}=\{0.22,0.1\}$ excluding 3BBF if the maximum tidal acceleration applied to the three bodies by the cluster during the encounter is >{${>}\{1\%,0.1\%\}$ of the three bodies’ maximum initial relative acceleration (see §A.1).
@@
Though we may increase certainty in the physical “isolation” of a 3BBF event by shrinking $\eta_{\rm H}$, what constitutes “isolated enough” is unconstrained in the $N$-body literature. A slew of $\eta_{H}$ models are generated to show the effect of asserting ever more conservative tidal isolation during 3BBF. A clear trend presents itself in Figure 3: smaller $\eta_{H}$ dramatically constrains the 3BB creation probability, especially for the widest binaries. This is a direct result of the 3BBF rate scaling as $\eta_{\rm H}^{3}$ (§ A.1). Depending on $\eta_{H}$, the number of 3BBs our model predicts to currently be within $200\,\rm{pc}$ ($1\,\rm{kpc}$) of the Sun varies from ${\sim}10^{5}$$10^{6}$ (${\sim}10^{6}$$10^{7}$).
@@
The choice of field evolution prescription heavily impacts the number and orbital element distributions of field binaries, with CS applying the most significant perturbations to individual binaries and the binary population as a whole. That said, both, the PM and CS binary field evolution prescriptions help bring the initially highly superthermal 3BB eccentricities in line with the more modestly superthermal observations. This is consistent with the conclusions of C. Hamilton (2022); C. Hamilton & S. Modak (2024), who postulated that the physical mechanism(s) responsible for creating the observed Gaia wide binaries would need to have a dramatic superthermal bias because of the thermalizing effect of stellar encounters and the Galactic tide after binary formation.
These 3BBs are the rare few that make a clean getaway from their natal cluster after formation. From Figure 1, the expected number of 3BBs escaping to the Galactic disk over a cluster’s life is
$N_{\rm bin,cl}\approx 0.6\left(\frac{\eta_{H}}{0.1}\right)^{3}\left(\frac{M_{\rm cl0}}{10^{4}M_{\odot}}\right)^{0.9},$ (3)
where $M_{\rm cl0}$ is the initial mass of the cluster, not the mass of the cluster at the point of 3BB emission. Despite the infrequency of 3BB formation, survival, and escape in individual low-mass clusters, our models suggest that the ubiquity of low-mass clusters makes them the largest contributors to a field 3BB population. The demographics of these binaries are discussed further in §§ 3.1 and 3.2, where the former strongly indicates 3BBF readily reproduces the observable Gaia superthermal wide binary population, alongside a meaningful contribution to the population of tight binaries.
@@ 3.1 Semi-major Axis and Eccentricity Distributions
To further compare the modeled and Gaia-observed binary populations, we express their respective SMA ($a$) and eccentricity ($\epsilon$) [^6] As a reminder to the reader, we employ $\epsilon$ as the eccentricity, not the eccentricity squared as is usually the case in the literature. distributions in Figures 3 and 5 in terms of the average projected separation (H.-C. Hwang et al., 2022b):
@@
Figure 3 provides the most visually compelling agreement, especially when we conservatively limit our maximum allowable 3BBF encounter volume to ${\sim}10\%$ of the local cluster Hill Radius ($\eta_{H}=0.1$). According to our model, the suppression in the wide binary population at higher $\langle s\rangle$ is directly attributable to binary dissolution by disk encounters (Equation C2). The 3BBF ${\sim}a^{2.5}$ proportionality for hard binaries, as expressed in D. C. Heggie (1975); J. Goodman & P. Hut (1993) and numerically reproduced in § A.2, correlates with the observed Gaia curve for separations ${<}10^{3}\,\rm{AU}$. However, there is substantial evidence that the tight‐binary regime is highly incomplete (K. El-Badry & H.-W. Rix, 2018), so this apparent agreement is likely coincidental, implying that 3BBF can only play a supporting role in directly populating tighter binaries.
The orbital perturbations by disk encounters on 3BBs is modeled with greater physical realism in the CS prescription. It does not merely consider the absolute destruction or conservation of binaries evolving under constant, isolated, secular evolution by the Galactic tides (as in PM), but enables stochastic binary evolution in SMA and eccentricity. A gentler fall-off in the separations of the wide 3BB population results, with most binaries with $a>10^{4}\rm AU$ born tighter, with $a$ between $10^{3}-10^{3.5}\rm AU$ (discussed further in § 3.2). These tighter binaries end up populating the widest subset of binaries through CS evolution, mitigating losses due to the rapid destruction of binaries with initial SMA $>10^{4}\rm AU$. The impact of CS on the eccentricity distribution is far stronger than that of PM.
@@
the $\alpha$ distribution is subthermal (less eccentric than thermal) if $\alpha<1$, thermal when $\alpha=1$, and superthermal if $\alpha>1$. It should be immediately apparent that the predicted and observed $\alpha$ values are not an exact match (variances within order unity). However, they are in good qualitative agreement; echoing exactly the observational findings of H.-C. Hwang et al. (2022b), our models show that binaries with $\langle s\rangle<10^{3}\rm{AU}$ tend subthermal while binaries with $\langle s\rangle>10^{3}\rm{AU}$ tend superthermal. The subthermal behavior is expected for binaries that form near the peak of the two-dimensional SMA–eccentricity distribution (where $a\approx R_{1}$, the radius of the three-body interaction volume; see § A.2).
The degree to which wide binaries exhibit a superthermal distribution varies dramatically between the PM and CS field evolution prescriptions (left panel and right panels of Figure 5, respectively). Because the PM prescription only has a cursory thermalizing effect, the eccentricity distributions of the widest binaries remain unreasonably superthermal. To the contrary, CS produces an $\alpha$ distribution in closer agreement with observations for wide binaries. As $\eta_{H}$ shrinks, both field evolution methods produce more eccentric binaries, though this is effect is small relative to how dramatically binary SMA and formation rate are affected by $\eta_{H}$.
@@
Figure 7: These five subfigures are SMA-focused, “1.5-dimensional” versions of the contour plots in Figure 6, where color (the extra “0.5th” dimension) distinguishes between different logarithmically-spaced bins ($i$) in SMA of width $\Delta a_{i}$. The bottom panel of each subfigure displays the cumulative density functions, $C(X,a_{i})$, of the integration variable $X$ specified on the panels’ horizontal axes, for each bin in SMA. The top panels of each subfigure display the corresponding probability density functions scaled by SMA bin width, $f(X)\Delta a_{\rm i}=\frac{\partial^{2}C(X,a_{\rm i})}{\partial X\partial a}\Delta a_{\rm i}$. This scaling enables a direct comparison of the relative contributions to the total G3R from each bin in SMA. The integration variables under examination are (a: $\{r/r_{\rm h}\}$) radial cluster location of 3BBF in terms of the natal cluster’s half-mass radius, (b: $\{\tilde{\tau}\}$) time between cluster formation and 3BBF as a fraction of the total cluster lifetime, (c: $\{M_{\rm cl}\}$) mass of the 3BB’s natal cluster at time $\tilde{\tau}$, (d: $\{R_{G}\}$) Galactocentric radius of natal cluster, (e: $\{t\}$) time the 3BB escapes its natal cluster, where $t=12\,\rm Gyrs$ is the present day. The widest binaries tend to be born later in time, both in $\tilde{\tau}$ and $t$, and further out, both in $r/r_{\rm h}$ and $R_{\rm G}$, than tight binaries. Across the entire parameter space, we predict that solar neighborhood 3BBs mostly formed in clusters with masses $M_{\rm cl}<10^{3}\,M_{\odot}$ (at $\tilde{\tau}$), with more than half born in $M_{\rm cl}<50M_{\odot}$ clusters.
@@
Figure 10: Contour plots (top) and an accompanying set of cumulative distribution functions (middle) illustrating the relationship between (i) the initial binary SMA, $a_{0}$, and its final SMA, $a$, following CS field evolution, and (ii) $a_{0}$ and the initial mass of the binary’s natal cluster, $M_{\rm cl0}$. In column (ii), we include the samples of binaries eventually disrupted through CS. The bottom panels show the fraction of disrupted field binaries given (i) $a_{0}$ or (ii) $M_{\rm cl0}$. The tighter the initial binary, the less its SMA tends to change. Critically, we predict that wide field 3BBs in the solar neighborhood today with SMA $a>4000\,\rm{AU}$ are overwhelmingly born with $a_{0}<4000\,\rm{AU}$. Hence, the wider the initial binary, the more rapidly it is destroyed in the field. Separately, the least massive clusters tend to produce the highest numbers of initially tight binaries, with over $50\%$ of $a_{0}<10^{3}\,\rm{AU}$ binaries born in clusters with $M_{\rm cl0}<100\,M_{\odot}$. Low-mass clusters contribute much of the initially widest binaries, with more than $50\%$ of the initially widest binaries emitted from clusters with $M_{\rm cl0}<500\,M_{\odot}$, but these binaries are overwhelmingly disrupted by CS.
@@ 3.2 Formation Characteristics of Solar Neighborhood Three-body Binaries
Figure 6 broadly highlights the underlying distribution and dimensional covariance found by evaluating the G3R. There are a few intuitive takeaways here: (i) most field 3BBs are created early in a cluster’s life when it is densest, according to the united equations of evolution (UEE, § A.4), with the smallest binaries born almost exclusively in a cluster’s infancy, (ii) very few binaries have orbital radii which migrate more than $3\,\rm{kpc}$ through the disk, as expected from the N. Frankel et al. (2020) diffusion model (§ B.1), and (iii) there is a sharp preference for 3BBs to have escaped clusters born at the onset of star formation in the thin disk ($t=6$ Gyrs in N. Frankel et al. (2020)).
@@
A more robust physical intuition may be found by examining “1.5-dimensional” versions of the contour plots relating binary SMA to cluster and MW history; they are displayed in Figure 7. Here, the dimensions (integration variables) of the G3R under scrutiny are (from left to right): $r/r_{h}$, the radial location within its host cluster where a 3BB is born, $\tilde{\tau}$, the fraction of the cluster’s lifespan elapsed when the 3BB forms and escapes, $M_{\rm cl}$, the mass of the cluster at the time of 3BBF and escape, $R_{G}$, the Galactocentric radius of the birth cluster’s circular orbit, and $t^{\prime}$, the age of the MW when the 3BB was born. Several interesting predictions are discernible here. First, nearly half of escaping 3BBs are born outside the host cluster’s Plummer core radius ($r\gtrsim 0.5r_{h}$), with wider binaries being born both further from the cluster’s center and later in the cluster’s life than tighter binaries. In fact, wide binaries tend to escape their birth cluster at a roughly uniform rate throughout the cluster’s life while ${>}50\%$ of the tightest 3BBs are born in the first ${\sim}10\%$ of the cluster’s life.
[…]
@@
It has long been postulated that dissolving clusters should be a reliable environment for dynamical wide binary formation (M. F. Sterzik & R. H. Durisen, 1998; N. Moeckel & M. R. Bate, 2010; M. B. N. Kouwenhoven et al., 2010; N. Moeckel & C. J. Clarke, 2011). Yet, we do not find isolated 3BBF to be a significant mechanism for generating solar neighborhood binaries at cluster dissolution $(\tilde{\tau}\approx 1)$, in either initially low- or high-mass clusters. Admittedly, our idealized treatment of star clusters as single-component Plummer models evolving under the UEE (§ A.4) is unlikely to reliably model cluster dissolution. The conditions for isolated 3BBF may also be rare in a cluster with few (${\lesssim}50$) bodies, though there is evidence that binary formation in strong interactions involving more than three bodies at the center of a cluster could be a significant source of tight binaries (A. Tanikawa, 2013). Such binaries are formed during particularly strong spikes in the cluster’s central density. Even our idealized cluster model exhibits density spikes at dissolution, so a more nuanced investigation of binary formation in high-density regions may be required to better assess whether dissolution is an evolutionary phase with significant binary generation. That said, the models of M. B. N. Kouwenhoven et al. (2010) found that extremely wide binaries predominantly form at the point of cluster dissolution, with a median binary SMA ${>}5000\,\rm{AU}$. Under CS (§ C.2), very few of these binaries can survive in the field.
@@
The left-hand panels of Figure 10 draw a relationship between the SMA before field evolution (pre-CS) and the final SMA of a 3BB population within the solar neighborhood today. As expected, the tighter the initial binary, the less CS shifts its SMA. A critical, population-level transition presents itself for the widest binaries: 3BBs with final SMAs ${>}4000\,\rm{AU}$ overwhelmingly escaped their natal clusters with initial SMAs ${<}4000\,\rm{AU}$. In other words, field evolution preferentially destroys the widest field 3BBs, replacing them with initially tighter binaries $({\sim}10^{3}$$10^{3.5}\,\rm{AU})$.
If we were to turn off disruption during CS field evolution, the distribution of wide binary SMA would become egregiously skewed to very wide SMA. The dearth of such very-wide binaries in our models is simply due to disruption during field evolution, and not 3BBF disfavoring very-wide binary production. Thus, any prospective wide binary formation channel that matches the observed Gaia SMA distribution, while neglecting field evolution, cannot account for most of the Gaia wide binary population. This implies that cluster dissolution may not be the primary source of solar neighborhood wide binaries, since it preferentially forms very-wide binaries before even considering field evolution and disruption.
@@
Finally, the relationship between initial binary SMA and the initial mass of the natal cluster is displayed in the right-hand panels of Figure 10. These distributions include the contribution of binaries which would ultimately be disrupted through CS, the left panel did not. Our models reinforce the general consensus that the field binary generation is dominated by low-mass clusters (N. Moeckel & C. J. Clarke, 2011).

### v3

@@ 3 Mock Three-body Binary Population versus Gaia Wide Binaries
Figure 1 displays the total number of solar neighborhood 3BBs we predict across all our models, alongside the wide binary $\left(\langle s\rangle>10^{3}\rm AU\right)$ tally. The “tidal isolation” parameter, $\eta_{H}$, is a nuisance parameter quantifying how isolated a three-body encounter is from tidal perturbations applied by its host cluster. As later discussed in § A.1, the interaction volume is defined to be
$R_{1}=\eta_{H}R_{H},$ (3)
where $R_{H}$ is the local Hill Radius. Shrinking $\eta_{H}$ further isolates three-body encounters, with $\eta_{H}=\{0.22,0.1\}$ excluding 3BBF if the maximum tidal acceleration applied to the three bodies by the cluster during the encounter is $\{1\%,0.1\%\}$ of the three bodies’ maximum initial relative acceleration.
@@
Though we may increase certainty in the physical “isolation” of a 3BBF event by shrinking $\eta_{\rm H}$, what constitutes “isolated enough” is unconstrained in the $N$-body literature. Several $\eta_{H}$ models are generated to show the effect of asserting ever more conservative tidal isolation during 3BBF. A clear trend presents itself in Figure 2: smaller $\eta_{H}$ dramatically constrains the 3BB creation probability, especially for the widest binaries. This is a direct result of the 3BBF rate scaling by ${\sim}\eta_{\rm H}^{3}$ (§ A.1). Depending on $\eta_{H}$, the number of 3BBs our model predicts to currently be within $100\,\rm{pc}$ ($1\,\rm{kpc}$) of the Sun varies from ${\sim}10^{4}$$10^{5}$ (${\sim}10^{6}$$10^{7}$).
@@
The choice of field evolution prescription heavily impacts the number and orbital element distributions of field binaries, with CS applying the most significant perturbations to individual binaries and the binary population as a whole. That said, both, the PM and CS binary field evolution prescriptions help bring the initial, highly superthermal 3BB eccentricities in line with the more modestly superthermal observations. This is consistent with the conclusions of C. Hamilton (2022); C. Hamilton & S. Modak (2024), who postulated that the physical mechanism(s) responsible for creating the observed Gaia wide binaries would need to have a dramatic superthermal bias because of the thermalizing effect stellar encounters and the Galactic tide impart following binary formation.
These 3BBs are the rare few that make a clean getaway—undergoing no encounters—from their natal cluster after formation. From Figure 1, the expected number of 3BBs escaping to the Galactic disk over a cluster’s life asymptotes to 0.20.20.50.5. The 3BBF creation and survival rates asymptote because the overwhelming majority of binaries that survive till dissolution are necessarily assembled late in the cluster’s life, when there are fewer than ${\sim}100$ bodies. Although three‑body binary formation (3BBF) events are intrinsically rare and the resulting binaries most often fail to survive within any single cluster, the sheer abundance of low‑mass clusters compensates for this scarcity. The demographics of these binaries are discussed further in §§ 3.1 and 3.2, where the former section explores 3BBF’s capacity to reproduce the observable Gaia superthermal wide binary population, and the latter examines the predicted solar-neighborhood binary population characteristics.
@@ 3.1 Semi-major Axis and Eccentricity Distributions
To further compare the modeled and Gaia-observed binary populations, we express their respective SMA ($a$) and eccentricity ($\epsilon$) [^5] As a reminder to the reader, we employ $\epsilon$ as the eccentricity, not the eccentricity squared as is usually the case in the literature. distributions in Figures 2 and 4 in terms of the average projected separation (H.-C. Hwang et al., 2022b):
@@
Figure 2 provides the most visually compelling agreement, especially when we conservatively limit our maximum allowable 3BBF encounter volume to ${\sim}10\%$ of the local cluster Hill Radius ($\eta_{H}=0.1$). According to our model, the suppression in the wide binary population at higher $\langle s\rangle$ is directly attributable to binary disruption by disk encounters (Equation C2). The 3BBF ${\sim}a^{2.5}$ proportionality for hard binaries, as expressed in D. C. Heggie (1975); J. Goodman & P. Hut (1993) and numerically reproduced in § A.2, mildly correlates with the observed Gaia curve for separations $10^{3}\,\rm{AU}$. However, there is substantial evidence that the tight‐binary regime is highly incomplete (K. El-Badry & H.-W. Rix, 2018), so any apparent agreement is most likely coincidental, implying that 3BBF can only play a supporting role in directly assembling tight binaries.
The orbital perturbations by disk encounters on 3BBs is modeled with greater physical realism in the CS prescription. It does not merely consider the absolute destruction or conservation of binaries evolving under constant, isolated, secular evolution by the Galactic tides (as in PM), but enables stochastic binary evolution in SMA and eccentricity. A gentler fall-off in the separations of the wide 3BB population results, with most binaries with final SMA $10^{4}\rm AU$ born tighter, with initial SMA between $10^{3}-10^{3.5}\rm AU$ (discussed further in § 3.2). These tighter binaries end up populating the widest subset of binaries through CS evolution, mitigating losses due to the rapid destruction of binaries with initial SMA $10^{4}\rm AU$.
@@
the $\alpha$ distribution is subthermal (less eccentric than thermal) if $\alpha<1$, thermal when $\alpha=1$, and superthermal if $\alpha>1$. It should be immediately apparent that the predicted and observed $\alpha$ values are not an exact match (variances within order unity), however they are in good qualitative agreement. Echoing the observational findings of H.-C. Hwang et al. (2022b), our models show that binaries with $\langle s\rangle<10^{3}\rm{AU}$ tend subthermal while binaries with $\langle s\rangle>10^{3}\rm{AU}$ tend superthermal. The subthermal behavior is expected for binaries that form near the peak of the two-dimensional SMA–eccentricity distribution (where $a\approx R_{1}$, the radius of the three-body interaction volume; see § A.2).
The degree to which wide binaries exhibit a superthermal distribution varies dramatically between the PM and CS field evolution prescriptions (left panel and right panels of Figure 4, respectively). Because the PM prescription only provides a cursory thermalizing effect, the eccentricity distributions of the widest binaries remain unreasonably superthermal. To the contrary, CS produces an $\alpha$ distribution in closer agreement with observations for wide binaries. As $\eta_{H}$ shrinks, both field evolution methods produce more eccentric binaries, though this is effect is small relative to how dramatically binary SMA and formation rates are affected by $\eta_{H}$.
@@
Figure 7: These five subfigures are SMA-focused, “1.5-dimensional” versions of the contour plots in Figure 6, detailing “where” and “when” 3BBs were assembled before migrating to the solar neighborhood. Color, the extra “0.5th” dimension, distinguishes between different logarithmically-spaced bins, $i$, in SMA of width $\Delta a_{i}$. The bottom panel of each subfigure displays the cumulative density functions, $C(X,a_{i})$, of the integration variable $X$ specified on the panels’ horizontal axes, for each bin in SMA. The top panels of each subfigure display the corresponding probability density functions scaled by SMA bin width, $f(X)\Delta a_{\rm i}=\frac{\partial^{2}C(X,a_{\rm i})}{\partial X\partial a}\Delta a_{\rm i}$. The integration variables under examination are (a: $\{r/r_{\rm h}\}$) radial cluster location of 3BBF in terms of the natal cluster’s half-mass radius, (b: $\{\tilde{\tau}\}$) time between cluster formation and 3BBF as a fraction of the total cluster lifetime, (c: $\{M_{\rm cl}\}$) mass of the 3BB’s natal cluster at time $\tilde{\tau}$, (d: $\{R_{G}\}$) Galactocentric radius of natal cluster, (e: $\{t^{\prime}\}$) time the 3BB’s natal cluster dissolves, where $t^{\prime}=12\,\rm Gyrs$ is the present day. The widest binaries tend to be born later in time, both in $\tilde{\tau}$ and $t^{\prime}$, and further out, both in $r/r_{\rm h}$ and $R_{\rm G}$, than tight binaries. Across the entire parameter space, we predict that solar neighborhood 3BBs mostly formed in clusters with masses $M_{\rm cl}(\tilde{\tau})<10^{2}\,M_{\odot}$, near the point of absolute dissolution of the cluster ($\tilde{\tau}{\sim}1$).
@@
Figure 10: Contour plots (top) and an accompanying set of cumulative distribution functions (middle) illustrating the relationship between (i) the initial binary SMA, $a_{0}$, and its final SMA, $a$, following CS field evolution, and (ii) $a_{0}$ and the initial mass of the binary’s natal cluster, $M_{\rm cl0}$. In column (ii), we include the samples of binaries eventually disrupted through CS. The bottom panels show the fraction of disrupted field binaries given (i) $a_{0}$ or (ii) $M_{\rm cl0}$. The tighter the initial binary, the less its SMA tends to change. Critically, we predict that wide field 3BBs in the solar neighborhood today with SMA $a>4000\,\rm{AU}$ are overwhelmingly born with $a_{0}<4000\,\rm{AU}$. Hence, the wider the initial binary, the more rapidly it is destroyed in the field. Separately, low-mass clusters tend to produce the highest numbers of binaries, with over $75\%$ of binaries born in clusters with $M_{\rm cl0}<10^{3}\,M_{\odot}$.
@@ 3.2 Formation Characteristics of Solar Neighborhood Three-body Binaries
Figure 6 broadly highlights the underlying distribution and dimensional covariance found by evaluating the G3R. There are a few intuitive takeaways here: (i) most field 3BBs are created early in a cluster’s life when it is densest, according to the united equations of evolution (UEE, § A.4), with the smallest binaries born almost exclusively in a cluster’s infancy, (ii) very few binaries have orbital radii which migrate more than $3\,\rm{kpc}$ through the disk, as expected from the N. Frankel et al. (2020) diffusion model (§ B.1), and (iii) there is a sharp preference for field 3BBs to be sourced from clusters born at the onset of star formation in the thin disk ($t=6$ Gyrs in N. Frankel et al. 2020).
@@
A more robust physical intuition may be found by examining “1.5-dimensional” versions of the contour plots relating binary SMA to cluster and MW history; they are displayed in Figure 7. Here, the dimensions (integration variables) of the G3R under scrutiny are (from left to right): $r/r_{h}$, the radial location within its host cluster where a 3BB is born, $\tilde{\tau}$, the fraction of the cluster’s lifespan elapsed when the 3BB forms and escapes, $M_{\rm cl}$, the mass of the cluster at the time of 3BBF and escape, $R_{G}$, the Galactocentric radius of the birth cluster’s circular orbit, and $t^{\prime}$, the age of the MW when the 3BB was born. Several interesting predictions are discernible here. First, roughly half of 3BBs are born outside the host cluster’s core ($r\gtrsim 0.5r_{h}$), with wider binaries being born both further from the cluster’s center and later in the cluster’s life than tighter binaries. Aside from the tightest 3BBs, most are born near dissolution in the final ${\sim}20\%$ of the cluster’s life.
The overwhelming majority of 3BBs originate from initially low-mass clusters ($M_{\rm cl0}<10^{3}\,M_{\odot}$) due to their dominance of the CIMF (Equation B4). Surviving wide binaries are born nearest to cluster dissolution, while tight binary formation occurs throughout a cluster’s life. Additionally, our model predicts that the tightest, solar neighborhood 3BBs should be born uniformly through the MW disk from $R_{G}=1$$8\,\rm{kpc}$ and early in MW star formation history, with $50\%$ born within the first $6\,\rm{Gyr}$ of the MW’s life (assuming a present MW age of $12\,\rm{Gyr}$). Wider binaries are born later in MW history and nearer to the solar neighborhood, an intuitive requirement due to the profound increase in a binary’s ionization cross section with an increasing binary SMA (§ C.1.1).
@@
Dissolving clusters have long been postulated to be reliable environments for dynamical wide binary formation (M. F. Sterzik & R. H. Durisen, 1998; N. Moeckel & M. R. Bate, 2010; M. B. N. Kouwenhoven et al., 2010; N. Moeckel & C. J. Clarke, 2011). We corroborate this, as we find isolated 3BBF to be a significant mechanism for generating solar neighborhood binaries when surviving till cluster dissolution $(\tilde{\tau}\approx 1)$. Admittedly, our idealized treatment of star clusters as single-component Plummer models evolving under the UEE (§ A.4) is unlikely to reliably model cluster dissolution. The conditions for isolated 3BBF may also be rare in a cluster with few (${\lesssim}50$) bodies, though there is evidence that binary formation in strong interactions involving more than three bodies at the center of a cluster could be a significant source of tight binaries (A. Tanikawa, 2013). Such binaries are formed during particularly strong spikes in the cluster’s central density. Even our idealized cluster model exhibits a density spike at dissolution, so a more nuanced investigation of binary formation near dissolution may be required to better assess the validity of our assumptions surrounding 3BBF at dissolution.
@@
The left-hand panels of Figure 10 draw a relationship between the SMA before field evolution (pre-CS) and the final SMA of a 3BB population within the solar neighborhood today. As expected, the tighter the initial binary, the less CS shifts its SMA. A critical, population-level transition presents itself for the widest binaries: 3BBs with final SMAs $4000\,\rm{AU}$ overwhelmingly escaped their natal clusters with initial SMAs $4000\,\rm{AU}$. In other words, field evolution preferentially destroys the widest field 3BBs, and widens initially tighter binaries $({\sim}10^{3}$$10^{3.5}\,\rm{AU})$. If we were to turn off disruption during CS field evolution, the distribution of wide binary SMA would become egregiously skewed to very wide SMA. The dearth of such very-wide binaries in our models is simply due to disruption during field evolution. Thus, any prospective wide binary formation channel that matches the observed Gaia SMA distribution, while neglecting field evolution, cannot account for most of the Gaia wide binary population.
@@
Finally, the relationship between initial binary SMA and the initial mass of the natal cluster is displayed in the right-hand panels of Figure 10. These distributions include the contribution of binaries which would ultimately be disrupted through CS, the left panel does not. Our models reinforce the general consensus that the field binary generation is dominated by low-mass clusters (N. Moeckel & C. J. Clarke, 2011).